How many ounces of a 20% saline solution must be added to 40 ounces of a 10% solution to produce a 16% solution?
Answer by stanbon(75887)

How many ounces of a 20% saline solution must be added to 40 ounces of a 10% solution to produce a 16% solution?
---------------------------------------
One-variable solution:
Equation:
0.20x + 0.10*40 = 0.16(x+40)
Multiply thru by 100 to get:
20x + 10*40 = 16x + 16*40
4x = 6*40
x = 6*10
x = 60 oz (amt of 20% solution needed in for the mixture)
=============================================================
Cheers,
Stan H.